OpenStudy (anonymous):

3. A textbook company keeps track of its daily cost for printing whole textbooks to different schools, but can only print up to 500 books a day. The daily cost (C) to print x number of textbooks is given by C(x) = 3.50x + 1200. 1. Using the function, determine one possible daily cost by evaluating the function for any input of your choice. 2. For what value of x does C(x) = 1900? 3. Interpret the statement C(20) = 1270. 4. Interpret the statement C(w) = p. 5. Explain whether C(40.5) = 1341.75 is correct and possible given the context of this problem.

OpenStudy (shreenithaa):

1. c(500)=3.50(500)+1200 ______________________________ 500 c=3.50+1200 c=1203.50 (FOR NUMBER 1, YOU COULD ACTUALLY PUT ANY NUMBER IN THERE FOR THE TEXTBOOKS)

OpenStudy (shreenithaa):

c(x)=1900 c(x)=3.50x+1200 original equation 3.50x+1200=1900 -1200 -1200 3.50x=700 / 3.50 x=700/3.50 c is x plugged in

OpenStudy (shreenithaa):

use the same idea for 3

OpenStudy (shreenithaa):

it says interpret so you have to guess for 4

OpenStudy (shreenithaa):

5. C(40.5) = 1341.75 c(40.5)=1341.75-(3.50(40.5)+1200) /////////////////////////////// 40.5 c=1341.75-(3.50+1200 c=1341.75-1203.5 c=138.25

OpenStudy (shreenithaa):

so i guess its correct

OpenStudy (shreenithaa):

is that enough or do u want me to do more?

OpenStudy (anonymous):

That's great. Thank you so much!

OpenStudy (shreenithaa):

ur welcome ;)

OpenStudy (shreenithaa):

what grade r u in?

OpenStudy (anonymous):

9th

OpenStudy (shreenithaa):

ooh im in 8th

OpenStudy (anonymous):

oh.. wow

OpenStudy (shreenithaa):

yeah maybe its because i have honors math and 9th graders go over it

OpenStudy (anonymous):

yeah probably

OpenStudy (shreenithaa):

k c ya

OpenStudy (anonymous):

thanks again bye